OpenStudy (anonymous):

6 men and 8 women are going to be assigned to a specific row of seats in a theater. If the 14 tickets for the numbered seats are given out at random, find the probability that exactly one womb is in one of the first 5 seats.

OpenStudy (anonymous):

woman not womb hahahahah

OpenStudy (anonymous):

no same thing

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

lololol XD

OpenStudy (anonymous):

does anyone know how to do this?

OpenStudy (anonymous):

This should be equivalent a number of combinations of n objects taken r at a time. Remember nCr = n! / [ (n - r)!r! a) There are 14C5 ways to select 5 people (to fill the first five seats) out of a total of 14 people (men and women). 14C5 =14!/[14-5]!5! =2002 b) There are 8C1 ways to select 1 woman out of 8. =8!/[8-1]!1! c) There are 6C4 ways to select 4 men out of 6 = 6!/[6-4]!4! d) There are 8C1*6C4 ways to select 1 woman out of 8 AND 4 men out of 6. 8C1*6C4= 120 ( from c and b above) Probability of choosing 1 woman and 4 men = P( 1 woman 4 Men) P(1 woman 4 Men) = 8C1*6C4/ 14C5 = 120/2002 =0.059940 Please verify accuracy of arithmetic. Better reading attachment

OpenStudy (anonymous):